# Flexible MOF Generation with Torsion-Aware Flow Matching

### v3

@@ 1 Introduction
Metal-organic frameworks (MOFs) are highly crystalline materials known for their permanent porosity, structural versatility, and broad applications in fields such as gas storage and separations [1, 2], catalysis [3], and drug delivery [4]. MOFs are designed by assembling chemical building blocks, i.e., metal clusters and organic linkers; their properties (e.g., pore size) can be tuned by swapping or modifying these components. This tunability has enabled the synthesis of over 100,000 distinct MOF structures to date [5], yet the theoretical space of possible MOFs is vastly larger, on the order of millions of structures [6]. Exploring this large design space is a grand challenge in materials science.
Computational modeling of MOFs, i.e., 3D structure prediction and generating new MOFs, poses significant challenges because of their structural complexity. Traditional methods, such as ab initio random search [7], attempt to design new MOFs or predict their structures by iteratively proposing combinations of building blocks and evaluating their stability using energy-based criteria. However, MOFs typically contain hundreds of atoms per unit cell, rendering these iterative approaches computationally prohibitive. Even state-of-the-art deep generative models developed for inorganic materials [8, 9] struggle to scale to the size and complexity of MOFs [10].
In response, researchers have proposed MOF-specific generative models such as MOFDiff [11] and MOFFlow [10]. While these methods represent significant progress, their design choices still constrain both chemical diversity and structural fidelity. Specifically, they (1) rely on a fixed set of predefined building blocks and (2) assume these building blocks are rigid. As a result, the design space is limited to recombinations of known components with fixed 3D conformations, hindering the discovery of novel chemistries and overlooking the intrinsic flexibility of organic linkers [12].
@@
MOFDiff [11]
MOFFlow [10] -
@@
In the first stage, an autoregressive Transformer generates a canonicalized sequence of MOF building blocks represented as SMILES [13], enabling the creation of entirely novel chemistries. These SMILES can be initialized to 3D structures using cheminformatics tools such as RDKit [14]. In the second stage, a flow-based model predicts the full 3D MOF structure by jointly modeling (1) rotations and translations of building blocks, (2) lattice structure, and (3) torsion angle for rotatable bonds in the organic linkers. Crucially, by explicitly modeling the torsion angle, MOFFlow-2 captures the conformational flexibility of organic linkers and no longer depends on predefined and rigid 3D conformations. See Table˜1 for comparison to prior works.
@@ 2 Related work
Generative models for MOFs. Unlike inorganic crystals, MOFs present unique challenges due to their large size, often containing hundreds of atoms per unit cell. MOFDiff [11] addresses this by generating coarse-grained structures via diffusion and decoding them using a fixed set of building blocks extracted from training data. Similarly, MOFFlow [10] assumes access to the 3D coordinates of building blocks and treats them as rigid bodies, learning block-level rotations and translations instead of predicting atom-level coordinates. In this work, we propose MOFFlow-2, a unified and flexible two-stage framework for both MOF generation and structure prediction. Unlike previous methods, MOFFlow-2 does not rely on a fixed building block library or require 3D coordinates as input, enabling greater chemical diversity and generality.
Equivariance and generative models. While SE(3) and E(3)-equivariant generative models have shown promising results on molecular tasks [15, 16, 17, 18], recent studies question whether equivariance is essential [19]. Notably, AlphaFold3 [20], Boltz-1 [21], and Proteína [22] achieve state-of-the-art results in modeling biomolecular complexes without using equivariant architectures. Similarly, Orb [23] and ADiT [24] employ non-equivariant neural networks to model interatomic potentials and generate molecule and crystal structures effectively. Following this trend, MOFFlow-2 adopts a non-equivariant Transformer architecture.
@@ 3.1 Structural representation of MOFs
MOF representation. The modular structure of MOFs allows a block-level representation based on their building blocks. For an MOF with $N$ atoms, we write $\mathcal{S}=(\mathcal{B}_{\text{3D}},\bm{q},\bm{\tau},\bm{\phi},\bm{\ell})$, where $\mathcal{B}_{\text{3D}}=\{\mathcal{C}^{(m)}\}_{m=1}^{M}$ denotes a set of $M$ building blocks. Each building block $\mathcal{C}^{(m)}=(A^{(m)},Y^{(m)})$ contains $N_{m}$ atoms with atom types $A^{(m)}\in\mathcal{A}^{N_{m}}$ and local coordinates $Y^{(m)}\in\mathbb{R}^{N_{m}\times 3}$. We represent the 3D structural degrees of freedom with rotations $\bm{q}=\{q^{(m)}\in SO(3)\}_{m=1}^{M}$, translations $\bm{\tau}=\{\tau^{(m)}\in\mathbb{R}^{3}\}_{m=1}^{M}$, and torsions $\bm{\phi}=\{\phi^{(m)}\in SO(2)^{P_{m}}\}_{m=1}^{M}$, where $P_{m}$ is the number of rotatable bonds in block $\mathcal{C}^{(m)}$. Following Jing et al. [16], we define a bond as rotatable if severing the bond creates two substructures, each with at least two atoms. The global coordinates $\bm{X}$ can be reconstructed by applying these transformations to local building block coordinates as $X^{(m)}=(q^{(m)},\tau^{(m)},\phi^{(m)})\cdot Y^{(m)}$ and concatenating them $\bm{X}=\operatorname{Concat}(X^{(1)},\dots,X^{(M)})$.
We also consider 2D molecular representations of MOF building blocks, denoted as $\mathcal{B}_{\text{2D}}=\{\mathsf{G}^{(m)}\}_{m=1}^{M}$. Each graph $\mathsf{G}^{(m)}=(\mathsf{V}^{(m)},\mathsf{E}^{(m)})$ consists of nodes representing atoms and edges corresponding to bonds in organic linkers. These bonds are inferred using OpenBabel [25] and RDKit [14] by analyzing the 3D structure $\mathcal{C}^{(m)}$ of the building block.
@@ 3.2 Flow matching on Riemannian manifolds
Flow matching (FM) trains a CNF by learning a vector field that transports samples from an arbitrary prior distribution $p_{0}$ to the target distribution $q$. Since MOFFlow-2 operates on $SO(2)$ and $SO(3)$, we here introduce flow matching generalized to Riemannian manifolds [26].
@@
While the vector field $u_{t}(x)$ is intractable, flow matching introduces a conditional formulation that yields a tractable and equivalent training objective [27]. Specifically, the probability paths $p_{t}(x)$ and vector field $u_{t}(x)$ can be expressed as marginalization over the conditional probability paths $p_{t}(x|x_{1})$ and conditional vector field $u_{t}(x|x_{1})$:
@@ 4 Autoregressive building block generation
In this section, we introduce our model for generating the MOF building blocks, which are used by the structure prediction model to generate the final 3D structure. Formally, we model the joint distribution as $p_{\theta}(\bm{q},\bm{\tau},\bm{\ell},\bm{\phi},\mathcal{B}_{\text{3D}})=p_{\theta}(\mathcal{B}_{\text{3D}})p_{\theta}(\bm{q},\bm{\tau},\bm{\ell},\bm{\phi}|\mathcal{B}_{\text{3D}})$ where $\mathcal{B}_{\text{3D}}$ denotes the 3D building block representations. To this end, we train a SMILES generative model $p_{\theta}(\mathcal{B}_{\text{2D}})$ to generate the 2D graph structure of each building block. Then we use pre-built metal building block library and RDKit [14] to obtain $\mathcal{B}_{\text{3D}}$, the initial 3D coordinates of the building blocks.
@@ 4.1 Autoregressive SMILES generation
The building block generator $p_{\theta}(\mathcal{B}_{\text{2D}})$ is an autoregressive model that generates MOF sequences, which contain metal clusters and organic linkers represented as SMILES [13]. This section outlines the definition of an MOF sequence, the training objective, and the implementation details.
@@
Next, Given the canonical MOF sequence $\mathcal{B}_{\text{2D}}$, we tokenize it into a sequence of discrete symbols $b_{1},\dots,b_{S}]$ by mapping substrings (e.g., atoms, bonds, or special tokens) into elements of a fixed vocabulary. We use a SMILES-aware regular expression [28] for tokenization, resulting in a vocabulary of size 59. See Section˜A.1 for an example of the tokenization process.
@@
Due to their complex chemistry, only a limited set of metal clusters is commonly used in MOFs [12]. Based on this observation, we construct a metal library of averaged 3D coordinates from the training data. Specifically, we group metal clusters by their canonical SMILES, select a random reference structure within each group, align all other structures to the reference using root mean square distance (RMSD) minimization, and compute the average of the aligned coordinates. [^1] Appendix B confirms this – our dataset has only 7–8 metal types with low RMSD variability.
In contrast, organic linkers exhibit greater structural diversity and flexibility [12], making the template-based approach impractical. Therefore, we use a cheminformatics toolkit (i.e., RDKit [14]) to initialize standard bond lengths and angles by optimizing the MMFF94 force field [29]. However, the resulting structures are only approximate, as the torsion angles around the rotatable bonds remain highly variable [16]. We therefore treat these torsion angles as learnable variables and model them with our structure prediction module to capture the full conformational flexibility of organic linkers.
@@ 5 MOF structure prediction with torsional degrees of freedom
Importantly, similar to MOFFlow [10], our structure prediction model MOFFlow-2 can be used independently of the building block generator. In fact, MOFFlow-2 is more general, as it can predict structures directly from the 2D building block $\mathcal{B}_{\text{2D}}$, without requiring known 3D conformations. This is more practical since the 3D structures are often unavailable or difficult to obtain in the real world.
@@ 5.1 Training algorithm
Prior distribution. We define the priors over each structural component as follows. Rotations and torsions are sampled independently from uniform distributions on their respective manifolds: $q\sim\mathcal{U}(SO(3))$ and $\phi\sim\mathcal{U}(SO(2))$. Translations are drawn from a standard normal distribution with the center of mass removed: $\bm{\tau}\sim\mathcal{N}(0,I_{M})$. For the lattice parameters, the lengths $(a,b,c)$ are sampled independently from log-normal distributions: $(a,b,c)\sim\operatorname{LogNormal}(\bm{\mu},\bm{\sigma})$, where $\bm{\mu}=(\mu_{a},\mu_{b},\mu_{c})$ and $\bm{\sigma}=(\sigma_{a},\sigma_{b},\sigma_{c})$ are estimated by maximum likelihood objective; the lattice angles $(\alpha,\beta,\gamma)$ are sampled uniformly from the range $60^{\circ},120^{\circ}] [$9, 10].
Conditional flows, vector fields, and flow matching loss. Following Chen and Lipman [26], we define conditional flows on each manifold as geodesic interpolations between a prior sample $\bm{z}_{0}$ and a target data $\bm{z}_{1}$, where $\bm{z}_{1}$ represents one of translation, rotation, torsion, or lattice parameter. Specifically, we define the flow as $\bm{z}_{t}=\exp_{\bm{z}_{0}}(t\log_{\bm{z}_{0}}(\bm{z}_{1}))$, where $t\in[0,1]$ and $\log(\cdot)$, $\exp(\cdot)$ denote logarithm map and exponential map on respective manifolds (see Section˜C.1 manifold-specific details). This allows parameterizing the vector field by $v_{t}(\bm{z}_{t};\theta)={\log_{\bm{z}_{t}}(\hat{\bm{z}}_{1})}/(1-t)$ where the neural network predicts the clean data $\hat{\bm{z}}_{1}$ from intermediate state $\bm{z}_{t}$. Furthermore, the conditional vector field is given by $u_{t}(\bm{z}_{t}|\bm{z}_{1})={\log_{\bm{z}_{t}}(\bm{z}_{1})}/(1-t)$.
@@ 5.2 Model architecture
Initialization module. The initial node embeddings $h_{i}^{(0)}\in\mathbb{R}^{D}$ for $i$-th atom is computed as $\smash{h_{i}^{(0)}=[E_{\theta}(a_{i}),f(a_{i}),x_{i},\bm{\ell}_{t},\varphi(t)]+\varphi(k)}$, where $E_{\theta}(a_{i})$ is a learnable atom type embedding, $f(a_{i})$ is a binary feature vector encoding atom-level features such as aromaticity [15], and $\varphi(\cdot)$ is sinusoidal embedding function, $t$ is timestep, and $k$ is the index of the building block that contains the $i$-th atom. The building block index $k$ resolves ambiguity between blocks with identical SMILES by lexicographic ordering similar to Section˜4.1. Full details are provided in Section˜D.1.
Interaction module. We adopt the Transformer encoder architecture for the 3D molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents atoms and $\mathcal{E}$ consists of edges between atoms within a specified cutoff threshold. For each edge $(i,j)\in\mathcal{E}$, we define the edge feature as $e_{ij}=[b(i,j),\operatorname{RBF}(\|x_{i}-x_{j}\|_{2})]$, where $b(i,j)\in\{0,1\}^{4}$ is a one-hot encoding of the bond type (single, double, triple, aromatic), or all zeros if no bond exists [15], and $\operatorname{RBF}(\cdot)$ is a radial basis function for interatomic distance.
@@
(Lattice parameter) To predict the lattice parameter, we mean-pool the node embeddings $\{h_{i}\}_{i=1}^{N}$ and apply a two-layer MLP with a GELU activation [30], i.e., set $\hat{\bm{\ell}}=\operatorname{MLP}(\frac{1}{N}\sum_{i=1}^{N}h_{i})$.
@@
(Rotation) Similar to translation, the rotation head applies block-wise attention pooling to node embeddings, followed by a two-layer MLP. Importantly, to represent rotations, we use a continuous 9D matrix representation based on singular value decomposition (SVD)[31, 32, 33]. We found this approach to be more stable than common parameterizations – such as Euler angles, axis-angle, or quaternions – which often lead to unstable training due to their discontinuities [31]. Concretely, during training, we supervise the predicted (possibly non-orthogonal) raw matrix output $\Delta\bm{Q}_{t\rightarrow{}1}$ using a mean squared loss against the target rotation $\Delta\bm{q}_{t\rightarrow{}1}$. During inference, our model projects $\Delta\bm{Q}_{t\rightarrow{}1}$ onto the nearest valid rotation matrix by solving the orthogonal Procrustes problem [31]:

### v4

@@ 1 Introduction
Metal-organic frameworks (MOFs) are highly crystalline materials known for their permanent porosity, structural versatility, and broad applications in fields such as gas storage and separations [li2018recent, qian2020mof], catalysis [lee2009metal], and drug delivery [horcajada2012metal]. MOFs are designed by assembling chemical building blocks, i.e., metal clusters and organic linkers; their properties (e.g., pore size) can be tuned by swapping or modifying these components. This tunability has enabled the synthesis of over 100,000 distinct MOF structures to date [moghadam2017development], yet the theoretical space of possible MOFs is vastly larger, on the order of millions of structures [wilmer2012large]. Exploring this large design space is a grand challenge in materials science.
Computational modeling of MOFs, i.e., 3D structure prediction and generating new MOFs, poses significant challenges because of their structural complexity. Traditional methods, such as ab initio random search [pickard2011ab], attempt to design new MOFs or predict their structures by iteratively proposing combinations of building blocks and evaluating their stability using energy-based criteria. However, MOFs typically contain hundreds of atoms per unit cell, rendering these iterative approaches computationally prohibitive. Even state-of-the-art deep generative models developed for inorganic materials [jiao2024crystal, millerflowmm] struggle to scale to the size and complexity of MOFs [kim2025mofflow].
In response, researchers have proposed MOF-specific generative models such as MOFDiff [fu2023mofdiff] and MOFFlow [kim2025mofflow]. While these methods represent significant progress, their design choices still constrain both chemical diversity and structural fidelity. Specifically, they (1) rely on a fixed set of predefined building blocks and (2) assume these building blocks are rigid. As a result, the design space is limited to recombinations of known components with fixed 3D conformations, hindering the discovery of novel chemistries and overlooking the intrinsic flexibility of organic linkers [park2024inverse].
@@
MOFDiff [fu2023mofdiff]
MOFFlow [kim2025mofflow] -
@@
In the first stage, an autoregressive Transformer generates a canonicalized sequence of MOF building blocks represented as SMILES [weininger1988smiles], enabling the creation of entirely novel chemistries. These SMILES can be initialized to 3D structures using cheminformatics tools such as RDKit [rdkit]. In the second stage, a flow-based model predicts the full 3D MOF structure by jointly modeling (1) rotations and translations of building blocks, (2) lattice structure, and (3) torsion angle for rotatable bonds in the organic linkers. Crucially, by explicitly modeling the torsion angle, MOFFlow-2 captures the conformational flexibility of organic linkers and no longer depends on predefined and rigid 3D conformations. See Table˜1 for comparison to prior works.
@@ 2 Related work
Generative models for MOFs. Unlike inorganic crystals, MOFs present unique challenges due to their large size, often containing hundreds of atoms per unit cell. MOFDiff [fu2023mofdiff] addresses this by generating coarse-grained structures via diffusion and decoding them using a fixed set of building blocks extracted from training data. Similarly, MOFFlow [kim2025mofflow] assumes access to the 3D coordinates of building blocks and treats them as rigid bodies, learning block-level rotations and translations instead of predicting atom-level coordinates. In this work, we propose MOFFlow-2, a unified and flexible two-stage framework for both MOF generation and structure prediction. Unlike previous methods, MOFFlow-2 does not rely on a fixed building block library or require 3D coordinates as input, enabling greater chemical diversity and generality.
Equivariance and generative models. While SE(3) and E(3)-equivariant generative models have shown promising results on molecular tasks [corsodiffdock, jing2022torsional, jumper2021highly, MatterGen2025], recent studies question whether equivariance is essential [wang2024swallowing]. Notably, AlphaFold3 [Abramson2024], Boltz-1 [wohlwend2024boltz1], and Proteína [geffner2025proteina] achieve state-of-the-art results in modeling biomolecular complexes without using equivariant architectures. Similarly, Orb [neumann2024orb] and ADiT [joshi2025allatom] employ non-equivariant neural networks to model interatomic potentials and generate molecule and crystal structures effectively. Following this trend, MOFFlow-2 adopts a non-equivariant Transformer architecture.
@@ 3.1 Structural representation of MOFs
MOF representation. The modular structure of MOFs allows a block-level representation based on their building blocks. For an MOF with $N$ atoms, we write $\mathcal{S}=(\mathcal{B}_{\text{3D}},\bm{q},\bm{\tau},\bm{\phi},\bm{\ell})$, where $\mathcal{B}_{\text{3D}}=\{\mathcal{C}^{(m)}\}_{m=1}^{M}$ denotes a set of $M$ building blocks. Each building block $\mathcal{C}^{(m)}=(A^{(m)},Y^{(m)})$ contains $N_{m}$ atoms with atom types $A^{(m)}\in\mathcal{A}^{N_{m}}$ and local coordinates $Y^{(m)}\in\mathbb{R}^{N_{m}\times 3}$. We represent the 3D structural degrees of freedom with rotations $\bm{q}=\{q^{(m)}\in SO(3)\}_{m=1}^{M}$, translations $\bm{\tau}=\{\tau^{(m)}\in\mathbb{R}^{3}\}_{m=1}^{M}$, and torsions $\bm{\phi}=\{\phi^{(m)}\in SO(2)^{P_{m}}\}_{m=1}^{M}$, where $P_{m}$ is the number of rotatable bonds in block $\mathcal{C}^{(m)}$. Following jing2022torsional, we define a bond as rotatable if severing the bond creates two substructures, each with at least two atoms. The global coordinates $\bm{X}$ can be reconstructed by applying these transformations to local building block coordinates as $X^{(m)}=(q^{(m)},\tau^{(m)},\phi^{(m)})\cdot Y^{(m)}$ and concatenating them $\bm{X}=\operatorname{Concat}(X^{(1)},\dots,X^{(M)})$.
We also consider 2D molecular representations of MOF building blocks, denoted as $\mathcal{B}_{\text{2D}}=\{\mathsf{G}^{(m)}\}_{m=1}^{M}$. Each graph $\mathsf{G}^{(m)}=(\mathsf{V}^{(m)},\mathsf{E}^{(m)})$ consists of nodes representing atoms and edges corresponding to bonds in organic linkers. These bonds are inferred using OpenBabel [o2011open] and RDKit [rdkit] by analyzing the 3D structure $\mathcal{C}^{(m)}$ of the building block.
@@ 3.2 Flow matching on Riemannian manifolds
Flow matching (FM) trains a CNF by learning a vector field that transports samples from an arbitrary prior distribution $p_{0}$ to the target distribution $q$. Since MOFFlow-2 operates on $SO(2)$ and $SO(3)$, we here introduce flow matching generalized to Riemannian manifolds [chen2024flow].
@@
While the vector field $u_{t}(x)$ is intractable, flow matching introduces a conditional formulation that yields a tractable and equivalent training objective [lipman2023flow]. Specifically, the probability paths $p_{t}(x)$ and vector field $u_{t}(x)$ can be expressed as marginalization over the conditional probability paths $p_{t}(x|x_{1})$ and conditional vector field $u_{t}(x|x_{1})$:
@@ 4 Autoregressive building block generation
In this section, we introduce our model for generating the MOF building blocks, which are used by the structure prediction model to generate the final 3D structure. Formally, we model the joint distribution as $p_{\theta}(\bm{q},\bm{\tau},\bm{\ell},\bm{\phi},\mathcal{B}_{\text{3D}})=p_{\theta}(\mathcal{B}_{\text{3D}})p_{\theta}(\bm{q},\bm{\tau},\bm{\ell},\bm{\phi}|\mathcal{B}_{\text{3D}})$ where $\mathcal{B}_{\text{3D}}$ denotes the 3D building block representations. To this end, we train a SMILES generative model $p_{\theta}(\mathcal{B}_{\text{2D}})$ to generate the 2D graph structure of each building block. Then we use pre-built metal building block library and RDKit [rdkit] to obtain $\mathcal{B}_{\text{3D}}$, the initial 3D coordinates of the building blocks.
@@ 4.1 Autoregressive SMILES generation
The building block generator $p_{\theta}(\mathcal{B}_{\text{2D}})$ is an autoregressive model that generates MOF sequences, which contain metal clusters and organic linkers represented as SMILES [weininger1988smiles]. This section outlines the definition of an MOF sequence, the training objective, and the implementation details.
@@
Next, Given the canonical MOF sequence $\mathcal{B}_{\text{2D}}$, we tokenize it into a sequence of discrete symbols $b_{1},\dots,b_{S}]$ by mapping substrings (e.g., atoms, bonds, or special tokens) into elements of a fixed vocabulary. We use a SMILES-aware regular expression [schwaller2019molecular] for tokenization, resulting in a vocabulary of size 59. See Section˜A.1 for an example of the tokenization process.
@@
Due to their complex chemistry, only a limited set of metal clusters is commonly used in MOFs [park2024inverse]. Based on this observation, we construct a metal library of averaged 3D coordinates from the training data. Specifically, we group metal clusters by their canonical SMILES, select a random reference structure within each group, align all other structures to the reference using root mean square distance (RMSD) minimization, and compute the average of the aligned coordinates. [^1] Appendix B confirms this – our dataset has only 7–8 metal types with low RMSD variability.
In contrast, organic linkers exhibit greater structural diversity and flexibility [park2024inverse], making the template-based approach impractical. Therefore, we use a cheminformatics toolkit (i.e., RDKit [rdkit]) to initialize standard bond lengths and angles by optimizing the MMFF94 force field [halgren1996merck]. However, the resulting structures are only approximate, as the torsion angles around the rotatable bonds remain highly variable [jing2022torsional]. We therefore treat these torsion angles as learnable variables and model them with our structure prediction module to capture the full conformational flexibility of organic linkers.
@@ 5 MOF structure prediction with torsional degrees of freedom
Importantly, similar to MOFFlow [kim2025mofflow], our structure prediction model MOFFlow-2 can be used independently of the building block generator. In fact, MOFFlow-2 is more general, as it can predict structures directly from the 2D building block $\mathcal{B}_{\text{2D}}$, without requiring known 3D conformations. This is more practical since the 3D structures are often unavailable or difficult to obtain in the real world.
@@ 5.1 Training algorithm
Prior distribution. We define the priors over each structural component as follows. Rotations and torsions are sampled independently from uniform distributions on their respective manifolds: $q\sim\mathcal{U}(SO(3))$ and $\phi\sim\mathcal{U}(SO(2))$. Translations are drawn from a standard normal distribution with the center of mass removed: $\bm{\tau}\sim\mathcal{N}(0,I_{M})$. For the lattice parameters, the lengths $(a,b,c)$ are sampled independently from log-normal distributions: $(a,b,c)\sim\operatorname{LogNormal}(\bm{\mu},\bm{\sigma})$, where $\bm{\mu}=(\mu_{a},\mu_{b},\mu_{c})$ and $\bm{\sigma}=(\sigma_{a},\sigma_{b},\sigma_{c})$ are estimated by maximum likelihood objective; the lattice angles $(\alpha,\beta,\gamma)$ are sampled uniformly from the range $60^{\circ},120^{\circ}] [$millerflowmm, kim2025mofflow].
Conditional flows, vector fields, and flow matching loss. Following chen2024flow, we define conditional flows on each manifold as geodesic interpolations between a prior sample $\bm{z}_{0}$ and a target data $\bm{z}_{1}$, where $\bm{z}_{1}$ represents one of translation, rotation, torsion, or lattice parameter. Specifically, we define the flow as $\bm{z}_{t}=\exp_{\bm{z}_{0}}(t\log_{\bm{z}_{0}}(\bm{z}_{1}))$, where $t\in[0,1]$ and $\log(\cdot)$, $\exp(\cdot)$ denote logarithm map and exponential map on respective manifolds (see Section˜C.1 manifold-specific details). This allows parameterizing the vector field by $v_{t}(\bm{z}_{t};\theta)={\log_{\bm{z}_{t}}(\hat{\bm{z}}_{1})}/(1-t)$ where the neural network predicts the clean data $\hat{\bm{z}}_{1}$ from intermediate state $\bm{z}_{t}$. Furthermore, the conditional vector field is given by $u_{t}(\bm{z}_{t}|\bm{z}_{1})={\log_{\bm{z}_{t}}(\bm{z}_{1})}/(1-t)$.
@@ 5.2 Model architecture
Initialization module. The initial node embeddings $h_{i}^{(0)}\in\mathbb{R}^{D}$ for $i$-th atom is computed as $\smash{h_{i}^{(0)}=[E_{\theta}(a_{i}),f(a_{i}),x_{i},\bm{\ell}_{t},\varphi(t)]+\varphi(k)}$, where $E_{\theta}(a_{i})$ is a learnable atom type embedding, $f(a_{i})$ is a binary feature vector encoding atom-level features such as aromaticity [corsodiffdock], and $\varphi(\cdot)$ is sinusoidal embedding function, $t$ is timestep, and $k$ is the index of the building block that contains the $i$-th atom. The building block index $k$ resolves ambiguity between blocks with identical SMILES by lexicographic ordering similar to Section˜4.1. Full details are provided in Section˜D.1.
Interaction module. We adopt the Transformer encoder architecture for the 3D molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents atoms and $\mathcal{E}$ consists of edges between atoms within a specified cutoff threshold. For each edge $(i,j)\in\mathcal{E}$, we define the edge feature as $e_{ij}=[b(i,j),\operatorname{RBF}(\|x_{i}-x_{j}\|_{2})]$, where $b(i,j)\in\{0,1\}^{4}$ is a one-hot encoding of the bond type (single, double, triple, aromatic), or all zeros if no bond exists [corsodiffdock], and $\operatorname{RBF}(\cdot)$ is a radial basis function for interatomic distance.
@@
(Lattice parameter) To predict the lattice parameter, we mean-pool the node embeddings $\{h_{i}\}_{i=1}^{N}$ and apply a two-layer MLP with a GELU activation [hendrycks2016gelu], i.e., set $\hat{\bm{\ell}}=\operatorname{MLP}(\frac{1}{N}\sum_{i=1}^{N}h_{i})$.
@@
(Rotation) Similar to translation, the rotation head applies block-wise attention pooling to node embeddings, followed by a two-layer MLP. Importantly, to represent rotations, we use a continuous 9D matrix representation based on singular value decomposition (SVD)[levinson2020analysis, bregier2021deep, geist2023rotations]. We found this approach to be more stable than common parameterizations – such as Euler angles, axis-angle, or quaternions – which often lead to unstable training due to their discontinuities [levinson2020analysis]. Concretely, during training, we supervise the predicted (possibly non-orthogonal) raw matrix output $\Delta\bm{Q}_{t\rightarrow{}1}$ using a mean squared loss against the target rotation $\Delta\bm{q}_{t\rightarrow{}1}$. During inference, our model projects $\Delta\bm{Q}_{t\rightarrow{}1}$ onto the nearest valid rotation matrix by solving the orthogonal Procrustes problem [levinson2020analysis]: